[14 marks] A cruise ship has docked in Fremantle harbour∗. Ofthe500peopleonboard,10have been diagnosed with the COVID-19 virus. The spread of the virus to other people on the ship is governed by the initial value problem dI dt= (1/500)I(500−I) with I(0) = 10 where I(t) is the number of people on board who have the virus and t is time (in days).

(a) Solve the initial value problem for I(t).

(b) Allowing for rounding, how long before everyone on the boat has the virus (in other words, solve I(t) = 499.5)?

(c) Use Euler’s method to compute a numerical solution for t ∈ [0,10] using a step size of ∆t = 0.5.

(d) Plot both of your solutions (from (a) and (c)) on the same set of axes and comment. ∗Please note, this is a completely made up question—it is for educational purposes only!
Answer #1

at - 500 $(0) I I(500-I I(0) = 10 dI 1/500-I) = dto Integrating on both sides, we get d1 Constant. ln (I) + ln (500- I - t +03 I (49tet) - sooet sooet 49 tet 1) Ilt-499.5 =+=? 499.5 = scoot 49+ et 244755 + 499.50 = 500et oset = 24475.5 et = 48951 t

n Tn f(Tn,In) In
0 0.0000 9.8000 10.0000
1 0.5000 14.4560 14.9000
2 1.0000 21.1487 22.1280
3 1.5000 30.5635 32.7023
4 2.0000 43.3791 47.9841
5 2.5000 59.9648 69.6736
6 3.0000 79.7934 99.6560
7 3.5000 100.6028 139.5527
8 4.0000 117.7649 189.8541
9 4.5000 124.9968 248.7366
10 5.0000 117.5006 311.2350
11 5.5000 96.2071 369.9853
12 6.0000 68.4923 418.0888
13 6.5000 43.1211 452.3350
14 7.0000 24.7416 473.8955
15 7.5000 13.3565 486.2663
16 8.0000 6.9559 492.9445
17 8.5000 3.5519 496.4225
18 9.0000 1.7951 498.1984
19 9.5000 0.9024 499.0960
20 10.0000 0.4524 499.5472

Analytic Data Euler Data 0.0000 2.0000 4.0000 6.0000 8.000010.00001 2.0000 tin days